\begin{thebibliography}
\small
\bibitem{HL} H. Hudzik and L. Maligranda, Some remarks on s -convex functions, Aequationes Mathematicae, 48.1 (1994), 100-111.
\bibitem{KBO} U. S. Kirmaci, M. K. Bakula, M. E. \"{O}zdemir and J. Pe\v{c}ari\'{c}, Hadamard-type inequalities for $s$-convex functions, Applied Mathematics and Computation, 193.1 (2007), 26-35.
\bibitem{AK} M. Avci, H. Kavurmaci and M. E. Ozdemir, New inequalities of Hermite-Hadamard type via $s$-convex functions in the second sense with applications, Applied Mathematics and Computation, 217 (2011), 5171-5176.
\bibitem{SYQ} Y. Shuang, H. P. Yin and F. Qi, Hermite-Hadamard type integral inequalities for geometric-arithmetically $s$-convex functions, Analysis, 33.2 (2013), 197-208.
\bibitem{KST} A. A. Kilbas, H. M. Srivastava and J. J. Trujillo, Theory and Applications of Fractional Differential Equations, North-Holland Mathematical Studies, Elsevier (North-Holland) Science Publishers, Amsterdam, London and New York, 2006.
\bibitem{GZS} F. Gao, W. P. Zhong and X. M. Shen, Applications of Yang-Fourier Transform to Local Fractional Equations with Local Fractional Derivative and Local Fractional Integral, Advanced Materials Research, 461 (2012), 306-310.
\bibitem{MSY} H. Mo, X. Sui and D. Yu, Generalized convex functions and some inequalities on fractal sets, Abstract and Applied Analysis, 2014.
\bibitem{M} H. Mo, Generalized Hermite-Hadamard Type Inequalities Involving Local Fractional Integrals, Abstract and Applied Analysis, 2014.
\bibitem{MS} H. Mo and X. Sui, Generalized s-convex function on fractal sets, Abstract and Applied Analysis, 2014.
\bibitem{Y} X. J. Yang, Advanced Local Fractional Calculus and Its Applications, World Science Publisher, New York, NY, USA, 2012.
\bibitem{WSZW} W. Wie, H. M. Srivastava, Y. Zhang, L. Wang, P. Shen and J.
Zhang, A Local Fractional Integral Inequality on Fractal Space Analogous to Anderson¡¯s Inequality, Abstract and Applied Analysis, 2014 (2014).
\bibitem{YBM} X. J. Yang, D. Baleanu and J. A. T. Machado, Mathematical aspects of heisenberg uncertainty principle within local fractional fourier analysis, Boundary Value Problems, 2013 (2013), 131-146.
\bibitem{GYK} Y. Zhao, D. F. Cheng and X. J. Yang, Approximation solutions for local fractional Schr\"{o}dinger Equation in the on-dimensional cantorian system, Advances in Mathematical Physics, 2013 (2013), 5.
\bibitem{YG1} A. M. Yang, Z. S. Chen, H. M. Srivastava and X. J. Yang, Application of the local fractional series expansion method and the variational iteration method to the Helmholtz equation involving local fractional derivative operators, Abstract and Applied Analysis, 2013 (2013), 6.




\end{thebibliography}

\documentclass[11pt,leqno]{article}
\usepackage{amsmath, amsthm, amsfonts, amssymb}
\usepackage{mathrsfs}
\newfam\msbfam
\font\tenmsb=msbm10    \textfont\msbfam=\tenmsb \font\sevenmsb=msbm7
\scriptfont\msbfam=\sevenmsb \font\fivemsb=msbm5
\scriptscriptfont\msbfam=\fivemsb

\newfam\bigfam
\font\tenbig=msbm10 scaled \magstep2   \textfont\bigfam=\tenbig
\font\sevenbig=msbm7 scaled \magstep2 \scriptfont\bigfam=\sevenbig
\font\fivebig=msbm5 scaled \magstep2
\scriptscriptfont\bigfam=\fivebig

\def\dint{\displaystyle\int}

\def\dfrac{\displaystyle\frac}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}{Lemma}[section]

\newtheorem{defn}{Definition}[section]
\newtheorem{pf}{Proof}

\begin{document}

{\begin{center}\bf\LARGE HERMITE-HADAMARD TYPE INEQUALITIES FOR GENERALIZED $s$-CONVEX FUNCTIONS ON REAL LINEAR FRACTAL SET  $\mathbb{R}^{\alpha}(0<\alpha<1)$

\end{center}}

\vspace{0.2cm}
 \begin{center}{ Huixia Mo\footnote[1]{
Correspondence: huixiamo@bupt.edu.cn\\
The research is supported by the NNSF of China (11471050, 11161042)} and Xin Sui

\vspace{0.2cm} School of Science, Beijing University of Posts and
Telecommunications, Beijing, 100876, China \vspace{0.2cm}\\}\end{center}

\vspace{0.3cm}

\begin{center}\begin{minipage}{5in}

\small {In the paper, we establish the Hermite-Hadamard type inequalities for the generalized s-convex functions in the second sense on  real linear fractal set $\mathbb{R}^{\alpha}(0<\alpha<1).$}\\

\vspace{0.3cm}
{\bf Key words}\;\;fractal sets, local fractional integral, Hermite-Hadamard type inequality, generalized s-convex function
\end{minipage}\end{center}

\begin{center}\section{Introduction}\end{center}

The convexity of functions is an important concept in the class mathematical analysis course.  In \cite{HL}, Hudzik and Maligranda considered two kinds of convex functions which are $s$-convex. And many important inequalities are established for the $s$-convex functions. For example, the Hermite-Hadamard¡¯s inequality is one of the best known results in the literature, see \cite{KBO,AK,SYQ} and so on.

In recent years, the fractal theory has received significantly remarkable attention\cite{KST}.
The calculus on fractal set can lead to better comprehension for the various real world models from science and engineering \cite{GZS}.

On the fractal set, Mo etal.\cite{MSY,M} introduced the definition of the generalized convex function and established Hermite-Hadamard type  inequality.
In \cite{MS}, the authors introduced two kinds of generalized $s$-convex functions on fractal sets $\mathbb{R}^{\alpha}(0<\alpha<1).$

 The definitions of the generalized $s$-convex functions are as follows:

\begin{defn}\cite{MS}\;Let $\mathbb{R_+}=[0,\infty].$ A function $f:\mathbb{R_+}\rightarrow\mathbb{R^{\alpha}}$
is said to be generalized $s$-convex $(0<s<1)$ in the first sense, if
$$f(\lambda_1u+\lambda_2v)\leq\lambda_1^{\alpha s}f(u)+\lambda_2^{\alpha s}f(v), \eqno (1.1)$$
for all $u,v\in\mathbb{R_+}$ and all $\lambda_1,\lambda_2\geq0$ with $\lambda_1^ s+\lambda_2^s=1.$ We denote this by $f\in GK_s^1.$
\end{defn}

\begin{defn}\cite{MS}\;A function $f:\mathbb{R_+}\rightarrow\mathbb{R^{\alpha}}$
is said to be generalized $s$-convex $(0<s<1)$ in the second sense, if
$$f(\lambda_1u+\lambda_2v)\leq\lambda_1^{\alpha s}f(u)+\lambda_2^{\alpha s}f(v), \eqno (1.2)$$
for all $u,v\in\mathbb{R_+}$ and all $\lambda_1,\lambda_2\geq0$ with $\lambda_1 +\lambda_2=1.$ We denote this by $f\in GK_s^2.$
\end{defn}

Note that, when $s=1$, the generalized $s$-convex function in both sense is generalized  convex function\cite{MS}.

Inspired by \cite{KBO,AK,M}, in the paper we will establish the Hermite-Hadamard type inequalities for generalized $s$-convex functions.

\section{Preliminaries}

Let us review the operations with real line number on fractal space  and
use the Gao-Yang- Kang's idea to describe the definitions of the
 local fractional derivative and local fractional integral \cite{Y,WSZW,YBM,GYK,YG1}.

If $a^\alpha,b^\alpha$ and $c^\alpha$ belong to the set $\mathbb{R}^\alpha$ of real line numbers , then\\
(1)\;\;$a^\alpha+b^\alpha$ and $a^\alpha b^\alpha$ belong to the set $\mathbb{R}^\alpha$;\\
(2)\;\;$a^\alpha+b^\alpha=b^\alpha+a^\alpha=(a+b)^\alpha=(b+a)^\alpha$;\\
(3)\;\;$a^\alpha+(b^\alpha+c^\alpha)=(a^\alpha+b^\alpha)+c^\alpha$;\\
(4)\;\;$a^\alpha b^\alpha=b^\alpha a^\alpha=(ab)^\alpha=(ba)^\alpha$;\\
(5)\;\;$a^\alpha(b^\alpha c^\alpha)=(a^\alpha b^\alpha)c^\alpha$;\\
(6)\;\;$a^\alpha(b^\alpha+c^\alpha)=a^\alpha b^\alpha+a^\alpha c^\alpha$;\\
(7)\;\;$a^\alpha+0^\alpha=0^\alpha+a^\alpha=a^\alpha,$ and $a^\alpha\cdot1^\alpha=1^\alpha\cdot a^\alpha=a^\alpha$.

\begin{defn}\cite{Y}\;\;Let $f:[a,b]\rightarrow \mathbb{R^\alpha}$ be a function such that
$$ |f(x)-f(y)|<c|x-y|^\alpha \;\;\;(x,y\in [a,b]),$$
for positive constants $c$ and $\alpha(0<\alpha\leq 1).$ This function is called H\"{o}lder continuous function. In this case, we think that $f(x)$ is  in the sapce $C_\alpha[a,b].$
\end{defn}

\begin{defn}\cite{Y} The local fractional derivative of $f(x)$ of order $\alpha$ at $x=x_0$ is defined by
$$f^{(\alpha)}(x_0)=\frac{d^\alpha f(x)}{dx^\alpha}|_{x=x_0}=\lim\limits_{x\rightarrow x_0}\frac{\triangle^\alpha(f(x)-f(x_0))}{(x-x_0)^\alpha},$$
where $\triangle^\alpha(f(x)-f(x_0))\cong\Gamma(1+a)(f(x)-f(x_0)).$
If there exists $f^{((k+1)\alpha)}(x)=\overbrace{D_x^{\alpha} \ldots D_x^\alpha}\limits^{k+1\; \mbox {times}}f(x)$ for any $x\in I\subseteq\mathbb{R},$  then we denoted
$f\in D_{(k+1)\alpha}(I)$, where $k=0,1,2\dots.$
\end{defn}

\begin{defn}\cite{Y}\;\;Let $f\in C_{\alpha}[a,b]$. Then the local fractional integral of the function $f(x)$ is defined by,
$$\begin{array}{cl}&_aI_b^{(\alpha)}f(x)\\
&
=\dfrac{1}{\Gamma(1+a)}\int_a^bf(t)(dt)^\alpha\\&
=\dfrac{1}{\Gamma(1+a)}\lim\limits_{\triangle t\rightarrow 0}\sum\limits_{j=0}^{N}f(t_j)(\bigtriangleup t_j)^\alpha,\end{array}$$
where $\triangle t_j=t_{j+1}-t_j,\;\triangle t=\max\{\triangle t_1,\triangle t_2,\triangle t_j,\ldots\}$ and $[t_j,t_j+1],$ $j=0,\ldots,N-1,$ $t_0=a,$ $t_N=b,$ is a partition of the interval $[a,b]$.
\end{defn}

\begin{lem}\cite{Y}\;\;Suppose that $g\in C_1[a,b]$ and  $f\in C_\alpha[g(a), g(b)],$ then
$$_{g(a)}I_{g(b)}^{(\alpha)}f(x)=_{a}I_{b}^{(\alpha)}f(g)(s)[g'(s)]^{\alpha}.$$
 \end{lem}

\begin{lem}\cite{Y}\\
(1)\;\; Suppose that $f(x)=g^{(\alpha)}(x)\in C_{\alpha}[a,b]$, then we have
   $$_{a}I_{b}^{(\alpha)}f(x)=g(b)-g(a).$$
(2)\;\;
 Suppose that $f(x), g(x)\in D_{\alpha}[a,b]$ and $f^{(\alpha)}(x), g^{(\alpha)}(x)\in C_{\alpha}[a,b]$, then we have
   $$_{a}I_{b}^{\alpha}f(x)g^{(\alpha)}(x)=f(x)g(x)\bigg|_{a}^{b}-_aI_{b}^{(\alpha)}f^{(\alpha)}(x)g(x).$$
 \end{lem}

\begin{lem}\cite{Y}\;\;\\
$$\dfrac{d^\alpha x^{k\alpha}}{dx^\alpha}=\dfrac{\Gamma(1+k\alpha)}{\Gamma(1+(k-1)\alpha)}x^{(k-1)\alpha}.$$

From the above formula and Lemma 2.2, we have
$$\dfrac{1}{\Gamma(1+\alpha)}\int_a^bx^{k\alpha}(dx)^{\alpha}=\dfrac{\Gamma(1+k\alpha)}{\Gamma(1+(k+1)\alpha)}(b^{(k+1)\alpha}-a^{(k+1)\alpha}),k\in R.$$
\end{lem}

\begin{lem}\cite{Y}\;\; \textbf{(Generalized H\"{o}lder¡¯s inequality)}
 Let $f,g\in C_{\alpha}[a,b]$, $p,q>1$ with $1/p+1/q=1$, then
 $$\dfrac{1}{\Gamma(1+\alpha)}\int_a^b|f(x)g(x)|(dx)^{\alpha}\leq\bigg(\dfrac{1}{\Gamma(1+\alpha)}\int_a^b|f(x)|^p(dx)^{\alpha}\bigg)^{1/p}\bigg(\dfrac{1}{\Gamma(1+\alpha)}\int_a^b|g(x)|^q(dx)^{\alpha}\bigg)^{1/q}.$$
\end{lem}

\begin{center}\section{Main Results}\end{center}

\begin{thm} Suppose that $f:\mathbb{R_+}\rightarrow\mathbb{R^{\alpha}}$ is a generalized $s$-convex function in the second sense, where $s\in(0,1).$ Let $a,b\in[0,\infty),$ $a<b.$ If $f\in C_{\alpha}[a,b]$, then the following inequalities hold:
$$\begin{array}{cl}\dfrac{2^{(s-1)\alpha}}{\Gamma(1+\alpha)}f\biggl(\dfrac{a+b}{2}\biggr)\leq
\dfrac{_{a}I_{b}^{\alpha}f(x)}{(b-a)^\alpha}\leq\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}\bigl(f(a)+f(b)\bigr).\end{array}\eqno{(3.1)}$$
\end{thm}

\begin{pf}
Let $x=a+b-t,$ then by Lemma 2.1, we have $_{\frac{a+b}{2}}I_b^{(\alpha)}f(x)={_a}I_{\frac{a+b}{2}}^{(\alpha)}f(a+b-t).$

 Moreover, $f$ is $s$ convex in the second sense, then
$$\begin{array}{cl}_aI_b^{(\alpha)}f(x)
&
={_a}I_{\frac{a+b}{2}}^{(\alpha)}\bigl[f(x)+f(a+b-x)\bigr]\\&
\geq 2^{\alpha s} {_a}I_{\frac{a+b}{2}}^{(\alpha)}f\biggl(\dfrac{a+b}{2}\biggr)\\&
=\dfrac{2^{(s-1)\alpha}}{\Gamma(1+\alpha)}(b-a)^\alpha f\biggl(\dfrac{a+b}{2}\biggr).\end{array}$$

In the other hand, let $x=b-(b-a)t$, $0\leq t\leq 1,$ then we get
$$\begin{array}{cl}_aI_b^{(\alpha)}f(x)&=(b-a)^\alpha {_0}I_1^{(\alpha)}f\bigl[ta+(1-t)b\bigr]\\&
\leq(b-a)^\alpha {_0}I_1^{(\alpha)}\bigl[t^{\alpha s}f(a)+(1-t)^{\alpha s}f(b)\bigr]\\&
=(b-a)\bigl[f(a) {_0}I_1^{(\alpha)}t^{\alpha s}+f(b) {_0}I_1^{(\alpha)}(1-t)^{\alpha s}\bigr].\end{array}$$

From Lemma 2.3, it is easy to see that
$${_0}I_1^{(\alpha)}t^{\alpha s}=\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)},$$
 and
 $${_0}I_1^{(\alpha)}(1-t)^{\alpha s}=\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}.$$

 So,
$$_aI_b^{(\alpha)}f(x)\leq(b-a)^\alpha\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}\bigl(f(a)+f(b)\bigr).$$

Therefore, we obtain
$$\dfrac{2^{(s-1)\alpha}}{\Gamma(1+\alpha)}f\biggl(\dfrac{a+b}{2}\biggr)\leq
\dfrac{_{a}I_{b}^{\alpha}f(x)}{(b-a)^\alpha}\leq\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}\bigl(f(a)+f(b)\bigr).$$
\end{pf}

\begin{thm}
Let $f: I\rightarrow \mathbb{R^{\alpha}}$, be a differentiable function on $I^0$($I^0$ is the interior of $I$) such that $f^{(\alpha)}\in C_{\alpha}[a,b]$, where $a,b\in I^0$, $a<b$. If $|f^{(\alpha)}|^q$ is generalized $s$-convex in the second sense on $[a,b]$ for some fixed $s\in(0,1)$ and $q\geq 1$, then
$$\begin{array}{cl}&\bigg|\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)\bigg|\\
&\leq\dfrac{(b-a)^\alpha}{2^\alpha}\bigg(\dfrac{\Gamma(1+\alpha)}{\Gamma(1+2\alpha)}\bigg)^{\frac{q-1}{q}}\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}
\biggl(\biggl(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg]^{\frac{1}{q}}\\
&\quad\times\bigg[\big|f^{(\alpha)}(a)\big|^q+\big|f^{(\alpha)}(b)\big|^q\bigg]^{\frac{1}{q}}\\
\end{array}$$

\end{thm}

To show Theorem 3.2  is right, we need the following Lemma.

\begin{lem}(\cite{M})\;\; Let $I\subseteq R$ be an interval, $f:I^{0}\subseteq R\rightarrow R^{\alpha}$  such that $f\in D_\alpha(I^0)$ and $f^{(\alpha)}\in C_\alpha[a,b]$ for $a,b\in I^0$ with $a<b$. Then the following equality holds:
$$\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)=\dfrac{(b-a)^\alpha}{2^\alpha}\dfrac{1}{\Gamma(1+\alpha)}\int_0^1(1-2t)^\alpha f^{(\alpha)}\big(ta+(1-t)b\big)(dt)^\alpha.$$
\end{lem}

Now, let us give the proof of Theorem 3.2.
\begin{pf}
 From Lemma 3.1, it is obvious that
$$\begin{array}{cl}
&\bigg|\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)\bigg|\\
\leq&\dfrac{(b-a)^\alpha}{2^\alpha}\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha \big|f^{(\alpha)}(ta+(1-t)b)\big|(dt)^\alpha.\end{array}\eqno{(3.2)}$$

Let us estimate $$\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha \big|f^{(\alpha)}(ta+(1-t)b)\big|(dt)^\alpha,$$ for $q=1$ and $q>1.$

{\bf Case I.}\;\; $q=1.$

 Since $|f^{(\alpha)}|$ is generalized $s$-convex on $[a,b]$ in the second sense, we can know that for any $t\in [0,1]$
$$\big|f^{(\alpha)}(ta+(1-t)b)\big|\leq t^{\alpha s}|f^{(\alpha)}(a)|+(1-t)^{\alpha s}|f^{(\alpha)}(b)|.$$

Then,  we have,
$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha \big|f^{(\alpha)}(ta+(1-t)b)\big|(dt)^\alpha\\
&\leq\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha \big[t^{\alpha s}|f^{(\alpha)}(a)|+(1-t)^{\alpha s}|f^{(\alpha)}(b)|\big](dt)^\alpha\\
&=\biggl[|f^{(\alpha)}(a)|\dfrac{1}{\Gamma(1+\alpha)}\int_0^1t^{\alpha s}|1-2t|^\alpha(dt)^\alpha+|f^{(\alpha)}(b)|\dfrac{1}{\Gamma(1+\alpha)}\int_0^1(1-t)^{\alpha s}|1-2t|^{\alpha}(dt)^\alpha\biggr].\\
\end{array}\eqno{(3.3)}$$

From Lemma 2.2 and lemma 2.3, it is easy to see that
$$\begin{array}{cl}
&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1t^{\alpha s}|1-2t|^\alpha(dt)^\alpha\\
&=\dfrac{1}{\Gamma(1+\alpha)}\bigg[\int_0^{\frac{1}{2}}t^{\alpha s}(1-2t)^\alpha(dt)^\alpha
+\dint_{\frac{1}{2}}^1t^{\alpha s}(2t-1)^\alpha(dt)^\alpha\bigg]\\
&=\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}\biggl(\bigg(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg].
\end{array}\eqno{(3.4)}$$
And, let $1-t=x,$ then by Lemma 2.1 and (3.4), we have
$$\begin{array}{cl}
&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1(1-t)^{\alpha s}|1-2t|^\alpha(dt)^\alpha\\
&=\dfrac{1}{\Gamma(1+\alpha)}\int_0^1 x^{\alpha s}|1-2x|^\alpha(dx)^\alpha\\
&=\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}\biggl(\bigg(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg].
\end{array}\eqno{(3.5)}$$

Thus, substituting (3.4) and (3.5) into (3.3), we have

$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha \big|f^{(\alpha)}(ta+(1-t)b)\big|(dt)^\alpha\\
&\leq\dfrac{\Gamma(1+\alpha)}{\Gamma(1+2\alpha)}
\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}
\biggl(\biggl(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg]\\
&\quad\times\bigg[\big|f^{(\alpha)}(a)\big|^q+\big|f^{(\alpha)}(b)\big|\bigg].\\
\end{array}\eqno{(3.6)}$$

Thus, from (3.2), we obtain

$$\begin{array}{cl}&\bigg|\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)\bigg|\\
&\leq\dfrac{(b-a)^\alpha}{2^\alpha}\dfrac{\Gamma(1+\alpha)}{\Gamma(1+2\alpha)}
\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}
\biggl(\biggl(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg]\\
&\quad\times\bigg[\big|f^{(\alpha)}(a)\big|^q+\big|f^{(\alpha)}(b)\big|\bigg].\\
\end{array}$$

{\bf Case II.}\;\; $q>1.$

Using the generalized H\"{o}lder's inequality(Lemma 2.4), we obtain
$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha\big|f^{(\alpha)}(ta+(1-t)b)\big|(dt)^\alpha\\&
=\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^{\alpha\frac{q-1}{q}}|1-2t|^{\alpha\frac{1}{q}}\big|f^{(\alpha)}(ta+(1-t)b)\big|(dt)^\alpha\\
&\leq\bigg(\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha(dt)^\alpha\bigg)^{\frac{q-1}{q}}\bigg(\dfrac{1}
{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha|f^{(\alpha)}(ta+(1-t)b)|^q(dt)^\alpha\bigg)^{\frac{1}{q}}.\end{array}\eqno{(3.7)}$$

It is obvious that
$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha(dt)^\alpha\\
&=\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)^\alpha(dt)^\alpha+
\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1(2t-1)^\alpha(dt)^\alpha=\dfrac{\Gamma(1+\alpha)}{\Gamma(1+2\alpha)}.\end{array}\eqno{(3.8)}$$

 Moreover, since $|f^{(\alpha)}|^q$ is generalized $s$-convex in the second sense on $[a,b],$ then
 $$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha\big|f^{(\alpha)}(ta+(1-t)b)\big|^q(dt)^\alpha\\
 &\leq\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha\bigg(t^{\alpha s}|f^{(\alpha)}(a)|^q+(1-t)^{\alpha s}|f^{(\alpha)}(b)|^q\bigg)(dt)^\alpha\\
&=|f^{(\alpha)}(a)|^q\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha t^{\alpha s}(dt)^\alpha+|f^{(\alpha)}(b)|^q\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha(1-t)^{\alpha s}(dt)^\alpha.\\
\end{array}$$

  From (3.3) and (3.4), it is easy to see that

$$\begin{array}{cl}
&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha t^{\alpha s}(dt)^\alpha\\
=&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha(1-t)^{\alpha s}(dt)^\alpha\\
=&\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}\biggl(\biggl(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg].
\end{array}$$

So,  $$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha\big|f^{(\alpha)}(ta+(1-t)b)\big|^q(dt)^\alpha\\
 =&\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}
\biggl(\biggl(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg]^{\frac{1}{q}}\bigg[\big|f^{(\alpha)}(a)\big|^q+\big|f^{(\alpha)}(b)\big|^q\bigg]^{\frac{1}{q}}.
\end{array}\eqno{(3.9)}$$

Thus, substituting (3.8) and (3.9) into (3.7), we have

$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha\big|f^{(\alpha)}(ta+(1-t)b)\big|(dt)^\alpha\\
\leq&\bigg(\dfrac{\Gamma(1+\alpha)}{\Gamma(1+2\alpha)}\bigg)^{\frac{q-1}{q}}\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}
\biggl(\biggl(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg]^{\frac{1}{q}}
\bigg[\big|f^{(\alpha)}(a)\big|^q+\big|f^{(\alpha)}(b)\big|^q\bigg]^{\frac{1}{q}}.\end{array}$$

So, from (3.2) it follows that
$$\begin{array}{cl}&\bigg|\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)\bigg|\\
&\leq\dfrac{(b-a)^\alpha}{2^\alpha}\bigg(\dfrac{\Gamma(1+\alpha)}{\Gamma(1+2\alpha)}\bigg)^{\frac{q-1}{q}}\bigg[\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}+\dfrac{\Gamma(1+\alpha)\Gamma(1+s\alpha)}{\Gamma(1+(s+2)\alpha)}
\biggl(\biggl(\dfrac{1}{2}\biggr)^{\alpha s}-2^\alpha\biggr)\bigg]^{\frac{1}{q}}\\
&\quad\times\bigg[\big|f^{(\alpha)}(a)\big|^q+\big|f^{(\alpha)}(b)\big|^q\bigg]^{\frac{1}{q}}.\\
\end{array}$$

Thus, we complete the proof of Theorem 3.2.
\end{pf}

\begin{thm}Let $f:$ $I\rightarrow \mathbb{R^{\alpha}},$ $I\subset[0,\infty)$, be a differentiable function on $I^0$ such that $f^{(\alpha)}\in C_{\alpha}[a,b]$, where $a,b\in I$ and $a<b$. If $|f^{(\alpha)}|^q$ is generalized $s$-convex in the second sense on $[a,b]$ for some fixed $s\in(0,1)$ and $q>1$, then
$$\begin{array}{cl}&\bigg|\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)\bigg|\\
&\leq\dfrac{(b-a)^\alpha}{2^\alpha}\bigg[\dfrac{\Gamma(1+\frac{q}{q-1}\alpha)}{2^\alpha\Gamma(1+\frac{2q-1}{q-1}\alpha)}\bigg]^{\frac{q-1}{q}}
\bigg(\dfrac{\Gamma(1+s\alpha)}{2^\alpha\Gamma(1+(s+1)\alpha)}\bigg)^{\frac{1}{q}}\\&
\quad\times\biggl[\biggl(|f^{(\alpha)}(a)|^q+\biggl|f^{(\alpha)}\biggl(\dfrac{a+b}{2}\biggr)\biggr|^q\biggr)^{\frac{1}{q}}+\biggl(\biggl|f^{(\alpha)}\biggl(\dfrac{a+b}{2}\biggr)\biggr|^q+|f^{(\alpha)}(b)|^q\biggr)^{\frac{1}{q}}\biggr].
\end{array}$$
\end{thm}

\begin{pf}
From Lemma 3.1, we have
$$\begin{array}{cl}&\bigg|\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)\bigg|\\
&\leq\dfrac{(b-a)^\alpha}{2^\alpha}\dfrac{1}{\Gamma(1+\alpha)}\int_0^1|1-2t|^\alpha |f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha\\
&\leq\dfrac{(b-a)^\alpha}{2^\alpha}\bigg[\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)|f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha\\
&\quad+\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1(2t-1)|f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha\bigg]
.\end{array}\eqno{(3.10)}$$

Let us estimate
$$\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)|f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha$$
and
$$\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1(2t-1)|f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha,$$
respectively.

 Using the generalized H\"{o}lder's inequality(Lemma 2.4), we obtain
$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)|f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha\\
&\leq\bigg(\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)^{\alpha\frac{q}{q-1}}(dt)^\alpha\bigg)^{\frac{q-1}{q}}
\bigg(\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}|f^{(\alpha)}(ta+(1-t)b)|^{\alpha q}(dt)^\alpha\bigg)^{\frac{1}{q}}.\end{array}\eqno{(3.11)}$$

It is easy to see that
$$\begin{array}{cl}\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)^{\alpha\frac{q}{q-1}}(dt)^\alpha=\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1(1-2t)^{\alpha\frac{q}{q-1}}(dt)^\alpha
=\frac{\Gamma(1+\frac{q}{q-1}\alpha)}{2^\alpha\Gamma(1+\frac{2q-1}{q-1}\alpha)}.\end{array}\eqno{(3.12)}$$

Let $|f^{(\alpha)}(ta+(1-t)b)|^q=U(t).$ Since $|f^{(\alpha)}|^q$ is a generalized $s$ convex function in the second sense , then $U(t)$ is  a generalized $s$ convex function in the second sense , too. Thus, from the right hand side of (3.1),
 it follows that

$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}|f^{(\alpha)}(ta+(1-t)b)|^q(dt)^\alpha \\ =&\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}U(t)(dt)^\alpha\\
\leq&\biggl(\dfrac{1}{2}-0\biggr)^\alpha\dfrac{\Gamma(1+s\alpha)}{\Gamma(1+(s+1)\alpha)}(U(0)+U(\frac{1}{2}))\\
=&\dfrac{\Gamma(1+s\alpha)}
  {2^\alpha\Gamma(1+(s+1)\alpha)}\biggl(\biggl|f^{(\alpha)}\biggl(\dfrac{a+b}{2}\biggr)\biggr|^q+|f^{(\alpha)}(b)|^q\biggr).\end{array}\eqno{(3.13)}$$

Thus, substituting (3.12) and (3.13) into (3.11), we get

$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)|f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha\\
&\leq\bigg(\dfrac{\Gamma(1+\frac{q}{q-1}\alpha)}{2^\alpha\Gamma(1+\frac{2q-1}{q-1}\alpha)}\bigg)^{\frac{q-1}{q}}
\bigg(\dfrac{\Gamma(1+s\alpha)}{2^\alpha\Gamma(1+(s+1)\alpha)}\bigg)^{\frac{1}{q}}
\biggl(|f^{(\alpha)}(a)|^q+\biggl|f^{(\alpha)}\biggl(\frac{a+b}{2}\biggr)\biggr|^q\biggr)^{\frac{1}{q}}.\end{array}\eqno{(3.14)}$$

Moreover,
$$\begin{array}{cl}\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1(2t-1)^{\alpha\frac{q}{q-1}}(dt)^\alpha=\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1(1-2t)^{\alpha\frac{q}{q-1}}(dt)^\alpha
=\frac{\Gamma(1+\frac{q}{q-1}\alpha)}{2^\alpha\Gamma(1+\frac{2q-1}{q-1}\alpha)}.\end{array}$$

And, similar to the estimate of (3.13), we have
   $$\begin{array}{cl}\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1|f^{(\alpha)}(ta+(1-t)b)|^q(dt)^\alpha\leq\dfrac{\Gamma(1+s\alpha)}
  {2^\alpha\Gamma(1+(s+1)\alpha)}\biggl(\biggl|f^{(\alpha)}\biggl(\dfrac{a+b}{2}\biggr)\biggr|^q+|f^{(\alpha)}(b)|^q\biggr).\end{array}$$

So, it is analogues to  the estimate of (3.11), we have

$$\begin{array}{cl}&\dfrac{1}{\Gamma(1+\alpha)}\int_0^{\frac{1}{2}}(1-2t)|f^{(\alpha)}(ta+(1-t)b)|(dt)^\alpha\\
&\leq\bigg(\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1(2t-1)^{\alpha\frac{q}{q-1}}(dt)^\alpha\bigg)^{\frac{q-1}{q}}
\bigg(\dfrac{1}{\Gamma(1+\alpha)}\int_{\frac{1}{2}}^1|f^{(\alpha)}(ta+(1-t)b)|^{q\alpha}(dt)^\alpha\bigg)^{\frac{1}{q}}\\
&\leq\bigg(\dfrac{\Gamma(1+\frac{q}{q-1}\alpha)}{2^\alpha\Gamma(1+\frac{2q-1}{q-1}\alpha)}\bigg)^{\frac{q-1}{q}}
\bigg(\dfrac{\Gamma(1+s\alpha)}{2^\alpha\Gamma(1+(s+1)\alpha)}\bigg)^{\frac{1}{q}}
\biggl(\biggl|f^{(\alpha)}\biggl(\frac{a+b}{2}\biggr)\biggr|^q+|f^{(\alpha)}(b)|^q\biggr)\biggr|^q\biggr)^{\frac{1}{q}}.\end{array}\eqno{(3.15)}$$

Thus, combining (3.10), (3.14) and (3.15), we obtain
$$\begin{array}{cl}&\bigg|\dfrac{f(a)+f(b)}{2^\alpha}-\dfrac{\Gamma(1+\alpha)}{(b-a)^\alpha}{_a}I_b^{(\alpha)}f(x)\bigg|\\
&\leq\dfrac{(b-a)^\alpha}{2^\alpha}\bigg[\dfrac{\Gamma(1+\frac{q}{q-1}\alpha)}{2^\alpha\Gamma(1+\frac{2q-1}{q-1}\alpha)}\bigg]^{\frac{q-1}{q}}
\bigg(\dfrac{\Gamma(1+s\alpha)}{2^\alpha\Gamma(1+(s+1)\alpha)}\bigg)^{\frac{1}{q}}\\&
\quad\times\biggl[\biggl(|f^{(\alpha)}(a)|^q+\biggl|f^{(\alpha)}\biggl(\dfrac{a+b}{2}\biggr)\biggr|^q\biggr)^{\frac{1}{q}}+\biggl(\biggl|f^{(\alpha)}\biggl(\dfrac{a+b}{2}\biggr)\biggr|^q+|f^{(\alpha)}(b)|^q\biggr)^{\frac{1}{q}}\biggr].
\end{array}$$
\end{pf}

Terefore, we complete the proof of Theorem 3.3.

\noindent{\bf Acknowledgments}\;\;{This work is supported by the National Natural Science Foundation of China (11161042, 11471050)}

\end{document}